If the perimeter of a rectangle is 44 cm and the area is 117 square cm, then what are the dimensions of the rectangle?
Answer by Alan3354(69443)

If the perimeter of a rectangle is 44 cm and the area is 117 square cm, then what are the dimensions of the rectangle?
-----------------
P = 2L + 2W = 44
L + W = 22
----
L*W = 117
Find a pair of factors of 117 that have a sum of 22.